UT 19 Jun 2019

Outline of tasks for night

  1. switch TCS to 2019A and let it run as long as possible (hopefully full night)
    • ssh tcs@tcs1
    • cd /lbt/tcs/2019A
    • ./tcs-cluster switch 2019A
    • (will probably hang; I will try to investigate this before switch)
    • restart GUIs,..
    • to revert, ./tcs-cluster switch 2018B (from /lbt/tcs/2019A)
  2. ARGOS/SOUL testing - highest priority for night.
    • SX-side only
      • Test AO Ref stars of different brightness (M92 R=9.4, R=15.1)
    • SX-ARGOS/SOUL + DX-ARGOS/FLAO test
      • Test AO Ref stars of different brightness (M92 R=9.4, R=15.1)
    • SX-ARGOS/SOUL + DX-ESM/FLAO test - 1 hour - target is M92 (K,H,J filters)
      • Test AO Ref stars of different brightness (M92 R=9.4, R=15.1)
    • SX-ARGOS/SOUL + DX-SL (M92 R=9.4, R=15.1)
  3. Collect more ESM (LUCI-2 DX) vs SL (LUCI-1 SX) - M92
    • R=9.4 AO Ref star
    • R=15.1 AO Ref star
  4. Isoplanetic test - N375 both sides - SX AO/SOUL - DX ESM/FLAO - M92 - compare degradation of corrections vs radius from AO Ref star
  5. N30 crowded field test - SX AO/SOUL - DX AO/FLAO - compare corrections across 30" FOV
  6. LUCI-1 Rotator ZP check (use AZ masks) -sky clear or thin clouds; seeing < 1.2"
  7. LUCI-1 AGW transform -sky clear or thin clouds; seeing < 1.2"
    • Decision tree: If ZP small - take confirmatory set, if ZP is large, update and take new transform set
    • Stone I or H 1st half
    • Stone L 2nd half
  8. LUCI-1 Photometric Zeropoint checks - requires clear skies (any seeing) - possible higher priority in order to address SX vs DX flux controversy
    • 2MASS calibrator fields - multiple visits over different airmasses, z,J,H,K,Ks filters

OSA: S. Allanson

ISp: J. Morris

ISA: B. Rothberg (Tucson)

AO: G. Taylor, D. Miller (both Tucson), J. Christou

Software: A. Coulson, P. Kubanek (both Tucson)

ARGOS: G. Rahmer, A. Cardwell (Tucson)

Mtn Manager: M. Wagner

Nightlog

UT 01:20 - started up LUCI GUIs, reconnected to TCS

UT 01:21 - initializing all on LUCI-1, initialize all but MOS on LUCI-2

UT 01:22 - taking test darks - 2.51 seconds

luci1/2.20190619.0001-0002

LUCI-2 shows Unknown in GUI - but should be ok - worked all through bino-imaging during last week

UT 02:08 - all done with the turbulence monitoring setup

UT 02:38 - enclosure open - waiting for sky to darken - will run P330E photometric standard observation Bino first

UT 02:49 - at 7th mag pointing star - waiting for sky to darken more

UT 03:05 - sending bino P330E phto standard script - Airmass 1.254-1.114 (K,H,J,Ks,z)

UT 03:13 - 1st offset failed left side - waiting for preset to complete - took 66 seconds - aborted wait for preset on telsvc

UT 03:14 - resending preset - will skip to 2nd dither position

UT 03:16 - moving to 2nd dither position

UT 03:19 - umac128 changed to 48 (defaulted back to 24 every time sub-system is restarted)

-BTW issuTrak wet down for a few minutes - just came back up (UT 03:24) - REALLY inconvenient!

-4D laser was left on during this observation

Files: luciN.20190618.XXXX

0003-0006 - K-band

0007-0010 - H-band

0011-0014 - J-band

0015-0018 - Ks-band

0019-0022 - z-band

(updated script to make K,H,J,Ks, 20x2.51 and keep z 26x2.51 (need all the photons we can get - this should shave a few minutes off the script)

UT 03:48 - completed star - reconfiguring for ARGOS-SX only


UT 04:06 - ready to move to NGC 5272 - doing a pointing check

UT 04:14 - sending collimation script

UT 04:16 - collimation script complete - sending preset - premature shot - Gustavo said "yes" but he really meant "no"

UT 04:21 - SH spots look bad, lasers seem off from nominal position

UT 04:27 - trying some beam sharpening - also had a closure for satellite and a bunch of airplanes flying overhead

UT 04:31 - putting a blind filter in while we wait...

UT 04:36 - resending preset - pause in after 1st row

UT 04:40 - waiting for IIF - loops are closed

UT 04:42 - skipping to next item didn't work

UT 04:43 - aborting and resending preset - waiting for satellite - abort wait for preset was highlighted again in IIF without me sending preset

UT 04:44 - cleared out the IIF, but after resending preset it failed - AO Ref Star missing?

UT 04:45 - resending preset again

UT 04:49 - loops opened - safe skips - paused LUCI script

UT 04:52 - gain changes are causing problems again

UT 04:56 - unable to reach Arcetri - no one on skype or zoom

UT 04:58 - shell RIPD

luci1.20190619.XXXX

0023 - images show elongation

0024-0025 - elongation continues - although fwhm pixels are 2.5-3

UT 05:03 - moving to M92 - higher elvation and to east to avoid wind - doing pointing check first

UT 05:06 - pointing good, sending collimation script

UT 05:08 - waiting for IIF on telsvc - hit abort wait for preset

UT 05:10 - sending ARGOS L1 M92 script - AORef = 9.4

UT 05:12 - images are elongated again

UT 05:16 - updated script so that its saved as a cube instead of integrated

UT 05:29 - putting in blind filter while we wait...

UT 05:31 - taking LUCI exposure - oops blind is in

UT 05:37 - taking another LUCI exposure (cube)

UT 05:38 - satellite closure

UT 05:54 - Lorenzo found the wrong config was being used - but now the lasers are off

UT 05:59 - starting M92 script AORef9.4 - integrated exposures

UT 06:02 - back to on-target - lost lasers - had brief satellite closure

UT 06:06 - NGS did not reacquire the AO Ref star, and lasers drifted

UT 06:07 - IIF says still waiting for offset to complete in telescope service, but IIFGUI shows offset is complete - Petr is checking on this

UT 06:11 - aborted wait for preset on telescope service - preset OK, clicking continue on LUCI GUI

UT 06:12 - this offset worked ok

UT 06:14 - FWHMs in 0035 ~2.3x2.5

UT 06:18 - 0.89" DIMM airmass 1.023

UT 06:28 - again - offset complete in IIF but LUCI does not see its complete - aborting wait for preset, skip to line I want and going - offset happy

UT 06:30 - switched to tip tilt only applied

UT 06:32 - 0044 exposure bad - shape on secondary needs to be corrected - shape of PSF flared and fell apart

UT 06:37 - AdSec RIP'd - recovering

UT 06:42 - resending preset - now back to K-band - tip-tilt only

UT 06:46 - now full ARGOS corrections

UT 06:51 - sending preset M92 (same PA, position center) for AORef = 15.2

UT 06:52 - arbitrator timeout - in closure window - aborted preset

UT 06:53 - resending preset - have 30min free (AORef = 15.2 M92)

UT 06:56 - IIF hungup again - abort wait for preset -taking exposure manually

UT 06:59 - continuing script - testing large offset and if we can reqacquire the faint AORef star

UT 07:03 - recovered from offset was OK

UT 07:08 - noticed in exposure 52-53 a severe elongation in FWHMs - LM issue?

UT 07:10 - paused LUCI, now beam sharpening

UT 07:17 - safe skips, AdSec went wonky, trying to recover

UT 07:20 - resuming LUCI script

UT 07:26 -noticed that with different AO Ref the shadowing is much worse. However, this same AO Ref was used for ESM, data reduced and did not show the same shadowing from the w-unit on LUCI-1

UT 07:30 - large offsets seems to cause problems when returning to the field

UT 07:31 - another IIF pause miscommunication - abort wait for preset on telsvc and skipping to desired location in script and going

UT 07:35 - flexure compensation loop on ARGOS is not closed and refusing to close

UT 07:38 - zombie processes?

UT 07:39 - lasers stopped propagating - software error

UT 07:43 - going to switch to doing P330E again

luci1.20190619.XXXX

K-band:

0026-0027 - elongated (savemode: integrated)

0028- left the blind in - ignore the exposure

0029 - stars elongated 45degrees

0030 - tip-tilt only image

0031 - test exposure after Lorenzo fixed the config (using wrong config) - FWHM 1.9-2.0 -20"+20" offset

0032- -20,+20 offset

0033 - -300,-300 offset (sky)

0034 - on-target - first 10seconds of exposure no laser

0035 - on target - loops closed

0036 - sky to -240"-240"

0037-0038 - on-target FWHM ~2pixels

0039 - offset to sky

H-band

0040-0041 - on target

0042- sky - FWHM ~0.67" (5.6pixels)

0043 - on target

0044 - tip/tilt only on lasers - image exploded - don't use - bad exposure

0045 - loops open - this is an H-band seeing limited

K-band:

0046 - tip-tilt only - FWHM 4.3 pixels

0047 - full ARGOS corrections - FWHM ~2.3pixels - look more like elongated than tip-tilt only

0048 - another FULL ARGOS correctoin (fixed LGS spots a little) - FWHM 1.8-2 pixels

0049-0050 - AORef star = 15.1 - on target - FWHM 2.7x2.9 pixels - some stars may seem elongated, get FWHMs 2.09x2.7 pixels - will have to look with better tools than RTD

0051 - off sky

0052-0053 - on-target - FWHMs elongated 1.9pix x 2.8pix

0054 - sky

0055- on target

0056-0057 on target - 1.9x2.7 pix FWHM - elongated

0058 - sky

H-band

0059 - on target - FWHMs 4.5pix-6pix - are we closed loop? What happened?

0060 - on target - FWHMs look elongated FWHMs 2.6x3.2pixels, 3.2pix x 3.6px

0061 - sky

0062 - on-target - stars very elongated 3.5x2.6pixel, 3pix x 2.3pixel


Photometric Standard Set #2

UT 07:44 - reauthorizing LUCI-LUCI

UT 07:46 - sending preset - Airmass 1.088-1.153

UT 07:50 - LUCI-2 instrument config thinks its still configuring but it looks to be correct - abort wait for preset is present on both side again - continuing script

UT 08:23 - Completed photometric standard

luciN.20190619.XXXX

L1 0063-0066 - K-band

0067-0070 - H-band

0071-0074 - J-band

0075-0078 - Ks-band

0079-0082 - z-band

L2 0023-0026 - K-band

0027-0030 - H-band

0031-0034 - J-band

0035-0038 - Ks-band

0039-0042 - z-band

UT 08:23 - plan now is ARGOS-binocular - M92

UT 08:27 - reconfiguring

UT 08:29 - dichroic problem SX (elevation 60)

UT 08:37 - arbirtrator SX failed, LGS controller down on SX

UT 08:47 - sending ARGOS-bino collimation script

UT 08:49 - wait for IIF issue again, abort for wait for preset and continue

UT 08:51 - loading script for Cube binocular M92 ARGOS

UT 08:55 - need to do beam sharpening

UT 09:00 - sending bino M92 preset

UT 09:01 - super slow on DX side

UT 09:02 - DX side hasn't done the acquisition yet

UT 09:05 - still waiting on RunAO on DX according to IIF

UT 09:08 - resending preset

UT 09:09 - DX still slower

UT 09:10 - resending preset for a 3rd time

UT 09:12 - preset completes on DX but doesn't proceed to runAO - hangs on DX

-skipping to next line to take data causes it to hang

UT 09:18 - sending preset again

UT 09:23 - LUCI-2 ARGOS only authorization - send a LUCI-2 script see if the same thing happens

UT 09:28 - sending preset LUCI-2 ARGOS only

-preset failed

-we are not propagating

UT 09:30 - sending again

UT 09:31 - same problem - calling Petr

-may be a 2019A ARGOS/FLAO issue as ARGOS/SOUL works ok and seeing limited binocular worked ok

UT 09:38 - reverting to 2018B TCS

UT 09:46 - going to test monocular DX ARGOS first then go back to binocular

UT 09:50 - sent collimation script - but AO was still running

UT 09:51 - collimaiton script L2 sent

UT 09:54 - sending ARGOS L2 script for M92 (AO Ref=9.4)

UT 09:56 - couldn't find star - may need to repoint - Steve nudged it into the field

UT 10:00 - resending preset

UT 10:05 - reconfiguring for binocular ARGOS

UT 10:11 - Dichroic movement failed on SX

UT 10:12 - sending collimation preset

UT 10:18 - collimation complete

UT 10:20 - seeing bubble 1.2-1.3" on DIMM

UT 10:26 - awaiting sharpening of spots on SX

UT 10:27 - resending preset (some time lost to acquiring and sharpening on SX even though preset had failed)

UT 10:29 - RunAO failed on SX - slope transmission is disabled?

UT 10:32 - says RunAO failed on SX - skipping to next line of script -but loops closed

UT 10:33 - PSF elongated on SX, nice and round on DX

UT 10:34 - DIMM 1.4"

UT 10:34 - DIMM 1.6"

tiptilt gain is 1.2 on SX and DX

UT 10:36 - Enrico suggests tiptilt gain should be much higher for SOUL compared to FLAO

-gains of 4 or 5 have worked (SX FLAO)

UT 10:41 - DIMM 1.6"-1.7"

UT 10:43 - resuming on-target

UT 10:45 - loop opened on SX

UT 10:47 - loops opened again on SX
UT 10:50 - givig up on ARGOS - will try to get one more set of data on P330E

luciN.20190619.XXXX

L1: 0083-0084 - on-target - PSF elongated

0085-0086 on target - loop opened

L2: 0043-0044 - M92 ARGOS K-band on-target - monocular (test)

0045-0046 - on target

0047-0048 - on target

UT 10:55 - AdSec set to 25 limit for PEPSI - taking advantage for a final swing at phot standard

-v3 of script will start with z,Ks,K,H,J

UT 10:58 - sending preset, copointing limit error again - need to reauthorize again to fix it

UT 11:00 - sending P330E preset again Airmass 2.227-2.3334

luciN.20190619.XXXX

L1: 0087-0090 z-band

L2: 0049-0052 z-band

UT 11:13 - closing up - will take internal flats for K,H,J,Ks,z

UT 11:23 - starting calibrations

Flats 5 lights off/5lights on
L1: 0091-0100 J-band

0101-0110 H-band

0111-0120 K-band

0121-0130 Ks-band

0131-0140 z-band

L2: 0053-0062 J-band

0063-0072 H-band

0073-0082 K-band

0083-0092 Ks-band

0093-0102 z-band

UT 11:55 - calibrations done - LUCIs put in safe config

-Steve noted a smell of smoke outside after we closed

-- BarryRothberg - 10 Jun 2019
Topic revision: r8 - 19 Jun 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback